• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name D. Lhuillier

    Log entry time 12:35:50 on June12,2004

    Entry number 125446


    keyword=Septum heating test

    Performed some tests with only Left septum on.
    Ramping up the beam current from 0 to 32uA seems to confirm the previous extrapolation from low current data.
    Raster size has no obvious effect on the septum heating
    Steering the beam at X=-2mm on target makes no difference either.

    FIGURE 1